OpenStudy (anonymous):

If f(x) = 2+{6}/{x}+{5}/{x^{2}}, find f'( x ). Find f'( 3 ). Find f''( x ). Find f''( 3 ).

OpenStudy (anonymous):

\[f(x) = 2+{6}/{x}+{5}/{x^{2}}, \]

OpenStudy (anonymous):

Ok well first you need to take the derivative. Do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright so have you worked that out yet?

OpenStudy (anonymous):

\[f \left( x \right)=2+6x ^{-1}+5x ^{-2}\] \[\frac{ d }{ dx }\left( x \right)^{n}=nx ^{n-1}\]

OpenStudy (anonymous):

i know how to get f'(x) and f'(3), but idk about f''(x) and f''(3)

OpenStudy (anonymous):

That second ' is just telling you to take the second derivative of the function. So if you have the first derivative worked out just take the derivative of whatever that is. I got -6-10x by the way, did you find the same thing?

OpenStudy (anonymous):

\[\frac{ d }{dx }f'\left( x \right)=f''\left( x \right) and so on.\]

OpenStudy (anonymous):

oh oops lol its (-6/x^2)-(10/x^3)

OpenStudy (anonymous):